OpenStudy (anonymous):

simplify the expression: (y+4)^3

OpenStudy (anonymous):



OpenStudy (anonymous):

(y+4)(y+4)(y+4) can u solve nw

OpenStudy (anonymous):

You can start like you did for the squaring problem, but then multiply that whole thing by another (y+4)

OpenStudy (anonymous):

oh! i see :) thanks again

OpenStudy (anonymous):

otherwise u can use the formula for cubes \[(a+b)^3=a^3+b^3+3ab(a+b)\]

OpenStudy (anonymous):

Just be sure when you do (y+4)( y^2+8y+16) to multiply/distribute the first y ALL the way across the second parenthesis AND to distribute the first 4 across all the way... you get a lot of terms, but then you can combine and simplify

OpenStudy (anonymous):

Yeah, that works too.. but best to know how to distribute if you forget the formula. Thanks @akash_809 I had forgotten about that...

OpenStudy (anonymous):

knowing formula always helps , you can find \[(a+b)^7=(a+b)^3 (a+b)^3(a+b)\] if u know the cubes

OpenStudy (anonymous):

Sure, knowing the formula is great, especially if you are doing a lot of these. It can be hard to remember some more complicated formulas if you aren't using them often... FOIL is easy, and the cube rule is pretty easy, but it never hurts to be able to re-derive a formula if you find that you can't remember something. Anyway, thanks again for the reminder :)

OpenStudy (anonymous):

ok @JakeV8